OpenStudy (anonymous):

sin2 θ = 2 sin θ + 3

OpenStudy (mathmate):

Hint: Domain of sin(x) is [-1,1] So [-1,1]=2[-1,1]+3 What is the value of \(sin(\theta)\) on the right hand side to satisfy the preceding equality? [-1,1] means the continuous interval from -1 to 1inclusively.

OpenStudy (perl):

is this the equation sin(2t ) = 2 sin(t) +3 2sin t cos t = 2 sin t + 3

OpenStudy (perl):

or is it sin^2(t) = 2 sin t + 3

OpenStudy (perl):

if its the latter, you can make it a quadratic expression sin^2(t) - 2sin (t) -3 = 0

OpenStudy (perl):

and then factor

OpenStudy (perl):

@mathmate the domain of sin x is (-oo,oo) , or all real numbers

OpenStudy (solomonzelman):

I think the space between the 2 and the theta that user posted should indicate that (s)he wants, \(\large\color{black}{ \sin^2(\theta) }\) \(\LARGE\color{white}{ \rm \left| \right| }\) So I would go with: \(\large\color{black}{ \sin^2(\theta) =2\sin(\theta)+3 }\) \(\LARGE\color{white}{ \rm \left| \right| }\)

OpenStudy (mathmate):

@perl oops, a moment of distraction! Thanks for heads up. :)